(4 pts) Evaluate the integral (5x – 10) dx using the net area of a region. (b) The positive and negative area are At = A =
Answer #1

Que S (5x-10).de 6 -6 6 (210) 6 8712) -6 = (5x2_102) (90-60)–(90+60) -120, 2 Now area= Negative S so dx dy X=-6 y-(52-10) 2 A9 AFE S (5x-10).de X=2 6 Ata (542 - 10x) 2 790-60) -(10-20) => A+= 30 + 10 =40 Answer